# HSM: Hierarchical Scene Motifs for Multi-Scale Indoor Scene Generation

### v2

@@ 5.3 Ablation
To evaluate the contribution of individual components, we ablate key elements of HSM: 1) w/o scene motif —  removing scene motifs and treating all objects as individual furniture pieces, 2) w/o scene spatial optimizer —  removing scene spatial optimization and directly using DFS solver positions to place scene motifs, and 3) w/o DFS solver —  removing DFS solver and directly using VLM-provided positions to place scene motifs. We assess the impact of these modifications on scene generation quality using SceneEval metrics. As shown in Tab. 4, removing any component results in worse performance than the full HSM. See supplement for a detailed analysis and qualitative comparison.
@@ 6 Conclusion
We presented Hierarchical Scene Motifs (HSM), a hierarchical framework for indoor scene generation that produces dense object arrangements across spatial scales in a 3D scene. Our approach models indoor scenes as a hierarchy of support regions, each to be populated with objects. By leveraging scene motifs, HSM generates object arrangements at multiple scales, from room-level furniture layouts to fine-grained small object placements, exploiting the recurring spatial patterns in indoor environments to maintain a consistent and structured approach. We believe HSM’s unified hierarchical framework represents a significant step toward generating densely populated and realistic indoor environments.
Acknowledgments
This work was funded in part by a CIFAR AI Chair, a Canada Research Chair, NSERC Discovery Grants, and enabled by support from the Digital Research Alliance of Canada. We also thank Jiayi Liu, Weikun Peng, and Qirui Wu for helpful discussions.
@@ Acknowledgments
We provide additional details about HSM’s motif library (Sec. A.1), support region extraction procedure (Sec. A.2), DFS solver formulation (Sec. A.3), scene spatial optimization details (Sec. A.4), our asset retrieval process (Sec. A.5), ablation analysis (Sec. A.6), computation cost (Sec. A.7), and limitations (Sec. A.8). We also provide details about the support region dataset we constructed (Appendix B), the user study instructions (Appendix C), and the VLM prompts used in HSM (Appendix D), along with extra scene-level qualitative examples (Fig. 9) and rendered results (Fig. 10).
@@ A.3 DFS Solver
$\sigma(p_{i1})=\alpha\left[1-\frac{\|p_{i1}-p_{i0}\|}{D_{\text{init}}}\right]_{\!+}+\delta_{\text{edge}}\beta\left[1-\frac{\phi_{s_{i},w}(p_{i1})}{D_{\text{wall}}}\right]_{\!+}$ (1)
@@
Figure 8: Qualitative comparison of ablation results.
@@
A.4 Scene Spatial Optimization
@@
When optimization is required, we create a combined mesh representation for the motif and optimize this representative object using a two-step iterative process: collision resolution followed by support validation. In the collision resolution phase, we first attempt a small vertical lift. If collisions remain, we apply a horizontal displacement away from the colliders. We use an adaptive step size based on penetration depth and apply movement constraints according to object types. Large objects remain floor-bound, wall objects maintain wall attachment, and ceiling objects preserve ceiling contact.
@@
A.5 Asset Retrieval
To generate scene motifs, HSM retrieves meshes from an object dataset based on the object category, CLIP similarity, and bounding box dimensions. During input description decomposition, we prompt the VLM to extract each object’s dimensions and style attributes. If these details are not explicitly provided, the VLM predicts them using common-sense knowledge. To retrieve the best-matching asset, we first filter the dataset by object category. We then use OpenCLIP [30] (ViT-H-14-378-quickgelu) to compute text-image similarity between the extracted descriptions and candidate objects in the dataset and select the top $k=5$ by similarity. Finally, the candidate whose dimensions best match those specified by the VLM is selected.
A.6 Ablation
We provide a detailed analysis of the ablation results in Tab. 4, examining how the removal of each component affects fidelity and plausibility metrics beyond the high-level comparison in the main paper. We also report w/o wall and ceiling —  removing wall and ceiling support regions in Tab. 6, and Fig. 8 shows a qualitative comparison.
@@ A.6 Ablation
w/o DFS solver. Without the DFS solver, scene motifs are directly placed by the VLM without enforcing geometric constraints such as wall alignment or staying within bounds. This leads to the largest drop in out of bound (OOR) among all ablations. On the plausibility side, collisions (COL) increase, support drops, and out of bounds (OOB) errors peak, as objects are often placed outside their intended regions or at invalid positions. The DFS solver is necessary to enforce layout constraints and ensure plausible object placement.
@@
A.7 Inference computational cost
The average cost for VLM calls per scene is approximately US\$0.50 and generation time is about 8 minutes (roughly US\$0.02 per object) with the current implementation. All experiments were run on a Intel i9-14900K CPU with 64GiB of RAM.
@@ A.8 Limitations and Future Work
Currently, our support region extraction algorithm assumes the geometry is curated with correct normals. Handling malformed mesh geometry is beyond our current scope, but we acknowledge this is an important future direction. We also note that the support region identification, through its geometric analysis, cannot distinguish accessible internal support regions (e.g., the inside of an open drawer) from inaccessible ones (e.g., a closed drawer). This can occasionally lead to unrealistic small object placements, such as placing a plant inside a closed drawer of a nightstand. A high number of iterations in HSM, with multiple VLM calls and optimization stages, can increase runtime. Lastly, although the current motif library is extensible; it could be improved by automatically learning new motifs from a VLM.
@@ Appendix B Support Region Dataset Details
To evaluate HSM’s capability in predicting support regions, we construct a dataset of 100 objects with annotated ground truth support regions. Specifically, we select 100 3D assets from the HSSD-200 dataset [34], spanning 12 categories: tables (36), TV stands (10), shelves (8), benches (6), sofas (6), desks (6), bathtubs (5), chairs (5), nightstands (5), sinks (5), trays (5), and racks (3). We specifically focus on two types of objects — those with multi-layer structures and those with irregularly shaped surfaces — due to their complex support regions, making them ideal for evaluation. For instance, TV stands, shelves, nightstands, trays, and racks typically feature multi-layer structures that require a non-trivial extraction process to identify valid support regions. Similarly, benches, sofas, bathtubs, chairs, sinks, and some tables and desks were selected for their curved or irregularly shaped surfaces, which pose additional challenges for support region prediction.
@@ 4. The amount of the piece (integer)
Figure 9: Extra qualitative comparisons at the scene level.
Figure 10: Rendered qualitative results.

### v3

@@ 5.3 Ablation
To evaluate the contribution of individual components, we ablate key elements of HSM: 1) w/o scene motif —  removing scene motifs and treating all objects as individual pieces, 2) w/o scene spatial optimizer —  removing scene spatial optimization and directly using DFS solver positions to place scene motifs, and 3) w/o DFS solver —  removing DFS solver and directly using VLM-provided positions to place scene motifs. Tab. 4 shows that removing any component results in worse performance. See Sec. A.5 for a detailed analysis and qualitative comparison.
@@ 6 Conclusion
We presented Hierarchical Scene Motifs (HSM), a hierarchical framework for indoor 3D scene generation that produces dense object arrangements across spatial scales. Our approach models indoor scenes as a hierarchy of support regions, each to be populated with objects. By leveraging scene motifs, HSM generates object arrangements at multiple scales, from room-level furniture layouts to fine-grained small object placements, exploiting recurring spatial patterns. We believe HSM’s unified hierarchical framework represents a significant step toward generating densely populated and realistic indoor environments.
Acknowledgments. This work was funded in part by the Sony Research Award Program, a CIFAR AI Chair, a Canada Research Chair, NSERC DG, and enabled by support from Digital Research Alliance. We thank Jiayi Liu, Weikun Peng, and Qirui Wu for helpful discussions.
@@
We provide additional details about HSM’s motif library (Sec. A.1), support region extraction procedure (Sec. A.2), DFS solver formulation (Sec. A.3.1), scene spatial optimization details (Sec. A.3.2), our asset retrieval process (Sec. A.4), ablation analysis (Sec. A.5), a breakdown of SceneEval results by difficulty (Tab. 5), computation cost and runtime analysis (Sec. A.6), open source VLM result (Sec. A.7), and limitations (Sec. A.8). We also provide details about the support region dataset we constructed (Appendix B), the user study instructions (Appendix C), and the VLM prompts used in HSM (Appendix D), along with extra scene-level qualitative examples (Fig. 10) and rendered scenes (Fig. 11).
@@ A.3.1 DFS Solver
$\sigma(p_{i1})=\alpha\left[1-\frac{|p_{i1}-p_{i0}|}{D_{\text{init}}}\right]_{+}+\delta_{\text{edge}}\,\beta\left[1-\frac{\phi_{s_{i},w}(p_{i1})}{D_{\text{wall}}}\right]_{+}$ (1)
@@
Figure 8: Qualitative comparison of ablation results. Removing individual components leads to noticeable artifacts, including colliding motifs, misaligned and colliding furniture, highlighting the role of each module in producing coherent and physically plausible scene layouts.
@@
A.3.2 Scene Spatial Optimization
@@
When optimization is required, we create a combined mesh representation for the motif and optimize this representative object using a two-step iterative process: collision resolution followed by support validation. In the collision resolution phase, we first attempt a small vertical lift. If collisions remain, we apply a horizontal displacement away from the colliders. We use an adaptive step size based on penetration depth and apply movement constraints according to its support region. Large objects remain floor-bound, wall objects maintain wall attachment, and ceiling objects preserve ceiling contact.
@@
A.4 Asset Retrieval
To generate scene motifs, HSM retrieves meshes from an object dataset based on the object category, CLIP similarity, and bounding box dimensions. During input description decomposition, we prompt the VLM to extract each object’s dimensions and style attributes. If these details are not explicitly provided, the VLM predicts them using common-sense knowledge. To retrieve the best-matching asset, we first filter the dataset by object category. We then use OpenCLIP [31] (ViT-H-14-378-quickgelu) to compute the text-image similarity between the extracted descriptions and candidate objects in the dataset and select the top $k=5$ by similarity. Finally, the candidate whose dimensions best match those specified by the VLM is selected.
A.5 Additional Quantitative Analysis
Performance by Difficulty. We report the breakdown of SceneEval evaluation results by difficulty (Easy, Medium, Hard) in Tab. 5. HSM maintains consistent performance across difficulty levels, whereas baseline methods (e.g., LayoutGPT, InstructScene) exhibit a significant drop in fidelity metrics (CNT, ATR) as prompt complexity increases.
Ablations. We provide a detailed analysis of the ablation results in Tab. 4, examining how the removal of each component affects fidelity and plausibility metrics beyond the high-level comparison in the main paper. We also report w/o wall and ceiling —  removing wall and ceiling support regions in Tab. 6, and Fig. 8 shows a qualitative comparison.
@@ A.5 Additional Quantitative Analysis
w/o DFS solver. Without the DFS solver, scene motifs are directly placed by the VLM without enforcing geometric constraints such as wall alignment or staying within bounds. This leads to the largest drop in out of bound (OOR) among all ablations. On the plausibility side, collisions (COL) increase, support (SUP) drops, and out of bounds (OOB) errors peak, as objects are often placed outside their intended regions or at invalid positions. The DFS solver is necessary to enforce layout constraints and ensure plausible object placement.
@@
A.6 Computational Cost and Runtime Analysis
All experiments present in the main paper were run on a Intel i9-14900K CPU with 64GiB of RAM. The average cost for VLM calls per scene is approximately US\$0.50 and generation time is about 8 minutes (roughly US\$0.02 per object) with the current implementation. For a representative scene that take 8 minutes and 37 seconds with 15 scene motifs, the total generation time averages to 35 seconds per scene motif. As shown in Fig. 9, the system’s speed is limited primarily by VLM inference latency rather than geometric computation, with scene motif generation as the main bottleneck.
Figure 9: Runtime breakdown. (Right) Total runtime distribution. (Left) Breakdown of the Scene Motif Generation stage.
A.7 Open Source VLM
To assess the accessibility and reproducibility of our framework, we evaluate HSM using the open-source InternVL3_5-30B-A3B [73] on the SceneEval [68] framework with the same set of descriptions. We use InternVL because initial tests with other open-source VLMs, including Qwen3-VL-30B-A3B-Instruct [4], were less stable. Qwen showed weaker instruction following and occasionally produced schema issues or hallucinated objects, which led to errors in mesh retrieval and generation. InternVL handled the pipeline instructions more consistently.
Specifically, HSM with InternVL records SceneEval Fidelity scores of 45.40% for object count (CNT), 47.81% for attribute (ATR), and 23.78% for object-object relationship (OOR), effectively matching or surpassing Holodeck (CNT 44.64%, ATR 39.42%, OOR 20.92%). In terms of physical plausibility, the open-source model maintains high validity with a support (SUP) score of 72.69% (vs. Holodeck’s 62.12%) and comparable collision rates (17.01% vs. 17.32%).
These results show that while fidelity scores using InternVL are lower than those achieved with GPT-4o (a significantly larger and more capable model), they still follow the same general trends. We observe that the open-source VLM exhibits reduced instruction-following capabilities, particularly during complex spatial reasoning tasks such as scene motif decomposition and inference. However, the fact that HSM with InternVL still achieves performance comparable to Holodeck (which uses GPT-4o) highlights the benefit of our hierarchical structure. This demonstrates that our framework is not overly dependent on any single proprietary model, offering a robust solution even when constrained to smaller, open-source model.
@@ A.8 Limitations and Future Work
Limitations. While HSM demonstrates strong performance in generating dense and realistic indoor scenes, several limitations remain. First, natural-language instructions often allow multiple plausible interpretations. In HSM, scene motifs operate locally within each subtree rather than as global structures, so the system may create different object layouts from the same description and cannot express relations that span across subtrees. For example, an instruction like “place the desk so it faces the bed” requires coordination between two subtrees that HSM does not handle. Second, the motif library is manually curated and contains a limited set of atomic spatial patterns. While hierarchical composition lets HSM express a wide range of arrangements, the system is still constrained by the coverage of these predefined primitives. As a result, it may struggle with more unusual or highly specific configurations that fall outside the library. Third, our support region extraction assumes well-curated 3D geometry with correct surface normals. Malformed meshes can produce incorrect support regions and result lead to misplaced objects. Forth, the current geometric analysis does not distinguish between accessible and enclosed support regions. This can cause small objects to be placed inside spaces that should be unreachable, such as placing a plant inside a cabinet). Finally, the iterative design of HSM requires multiple VLM queries and optimization steps, which increases runtime, especially for larger scenes.
Future Work. There are several promising directions for extending HSM. Incorporating a mechanism to distinguish accessible from enclosed support regions would improve placement realism. Runtime efficiency could be improved through parallel processing or using GPU acceleration. In addition, automatically learning new motif types from VLMs or existing scene datasets would enable the system to expand its library dynamically and support more diverse indoor layouts.
@@ Appendix B Support Region Dataset Details
To evaluate HSM’s capability in predicting support regions, we construct a dataset of 100 objects with annotated ground truth support regions.
Specifically, we select 100 3D assets from the HSSD-200 dataset [35], spanning 12 categories: tables (36), TV stands (10), shelves (8), benches (6), sofas (6), desks (6), bathtubs (5), chairs (5), nightstands (5), sinks (5), trays (5), and racks (3). We specifically focus on two types of objects — those with multi-layer structures and those with irregularly shaped surfaces — due to their complex support regions, making them ideal for evaluation. For instance, TV stands, shelves, nightstands, trays, and racks typically feature multi-layer structures that require a non-trivial extraction process to identify valid support regions. Similarly, benches, sofas, bathtubs, chairs, sinks, and some tables and desks were selected for their curved or irregularly shaped surfaces, which pose additional challenges for support region prediction.
@@ 4. The amount of the piece (integer)
Figure 10: Extra qualitative comparisons at the scene level. Among all methods, HSM produces the most consistent room layouts and object arrangements with respect to the input descriptions.
Figure 11: Rendered qualitative results. HSM is able to generate realistic and densely populated scenes. The scenes also contain smaller objects and are aligned with user input.